Libby Area Technical Assistance Group, Inc.
PO Box 53 • Libby, MT 59923
July 12th, 2011 • Regular Board Meeting Minutes
(Note: bold items within paragraphs are motions made and voted on as well as action items agreed upon.)
Board Member Attendees:
- Eileen Carney
- JoElyn Brus
- Mike NobleKaren Horton
- Phillip Erquiaga
- Donna Martin
- Tracy McNew, Grant Administrator
- Terry Spear, Technical Advisor (via phone)
- Jennifer Berig, EPA
- John Podolinsky, MT DEQ
- Dania Zinner, EPA
- Mike Cirian, EPA
- Mel Parker
- Liz Fagen, EPA
- Carolyn Rutland, MT DEQ
- Rebecca Thomas, EPA
- Estelle Shiroma, Tech Law
- David Berry, EPA
- Christina Progess, EPA
- Monique Ammidown, Summit
- Josh Haynes, Summit
- Linda Gaines, EPA HQ
- Alan Gerstenecker
- Thomas Cook, CDM
- Steve Wharton, EPA
1. Meeting began at approximately 6:00pm.
2. Introductions were made around the room.
3. Special Presentations
a) Rebecca Thomas, EPA update re: OU1 Riverfront Park. (A handout was provided)
· The City of Libby submitted a design which the EPA has agreed to design their remedy around.
· There will be at least 18 inches of clean cover at the site.
· Some work will not be completed this year due to high water levels. E.g. boat ramp
· Final design is expected next week with construction starting this summer.
· A post construction risk assessment will be done once remediation is complete
· EPA is working with the city on an irrigation system for the park
· Mike N. asked if the road would be asphalt. Rebecca answered that it will be gravel as far as she knows.
· Phillip E. asked what will happen along the riverbank. Rebecca and Mike C. answered that the soils will be excavated and the bank will be armored. It is designed to be a permanent fix and future flood conditions are planned for.
· Mel P. asked if the old riprap would be removed. Rebecca answered that they will add to the existing riprap to extend below the low water mark and above the high water mark. Some vegetation will be mixed in with the riprap.
· Mike N. asked about the water loading station. Rebecca and Mike C. answered that it will remain but may be changed by the City.
b) Mike Cirian, EPA presentation on cleanup process. (PowerPoint)
· Mike described the superfund site, the operable units and their boundaries which have expanded over time. In 2002 the Libby superfund site encompassed 180 square miles.
· The structure of Libby’s amphibole fiber is straight and needle-like vs. curly like Chrysotile.
· Current triggers for clean-up were described including visible vermiculite and gravel driveways.
· Vermiculite is being left behind in walls and sealed off. ERS, Environmental Resource Specialist’s should be called if any work is being done to structures.
· Clean-up started with worst first approach, but now most properties have approximately the same level of contamination.
· Families and pets are relocated during remediation.
· Work is done with wet soils and workers wearing respirators as a precaution.
· Average level of soil removal is 12 inches but can go to 3 feet if needed. At three feet if contamination remained below, fencing would be laid, but this is uncommon.
· Safe transport of contaminated materials was discussed including using sealed end gates on trucks.
· Building materials go to the landfill and contaminated soils go to the mine site for disposal.
· It was discussed why the mine was not made in to a landfill. Mike C. suggested that the landfill offers closer hauling, single handling and daily cover.
· No remedy is in place for the mine site at this time.
· Terry S. asked if personal air monitors are used by all workers. Mike C. answered no. There is sampling going on everyday though.
· Long term outlook and the need to always have an ERS type program were discussed.
· Burning of materials was discussed and Kristina P. states that a test burn is being planned for a burn chamber in Research Triangle Park in NC. Dave B. explained that data collected from such a burn could relate particle size to distribution and allow more understanding of the danger fibers may pose in a forest fire.
c) Technical Advisor Report –Terry Spear
· Terry had questions regarding the OU2 Draft RA report, but Rebecca T. requested he put them in writing for submission.
4. DEQ Report- John Popolinsky
· DEQ is working with TetraTec to do property evaluation studies and ABS in Troy.
· A question was asked about role clarification between EPA and DEQ at this superfund site. It was clarified that:
o EPA is the lead on the project in Libby; DEQ is the lead on the project in Troy.
o The two agencies work in conjunction, but when EPA leaves, DEQ will be responsible for O&M.
o DEQ has direct responsibility to the people of MT. They act as a partner to EPA but also as a protector of MT citizens.
5. W.R. Grace Report (No Report) -
6. Meeting minutes from May and June of 2011 were reviewed and unanimously approved.
7. Treasurer’s Report –JoEllen Brus
- After reimbursement 87 and 88 were received the balance is $37,680.86.
8. Sub-Committee Reports
a) Communications Committee- Phillip Erquiaga (Chairperson) – No Report
b) Nomination Committee- LeRoy Thom (Chairperson) – No Report
c) O&M Attendance Rotation Committee- Mike N./LeRoy T. (Co-chairs) – No Report
d) Technical Advisor Committee- Dr. Black (Chairperson) – No Report
9. Old Business
a) Budget updated- Tracy will check with Linda A. to make sure she received it.
b) Newsletter- Phillip will complete this.
c) Dr. Black’s nomination to EPA review board & letters to senators. – He received letter from EPA.
Tracy will follow-up to see if he would like to pursue this. It as unanimously approved that
contingent on his decision TAG submits letters to both senators and the EPA regarding this
d) CAG/ TAG Symposium- LeRoy to follow-up with presentations to City and County.
10. New Business
a) It was unanimously approved to renew the Web Master contract for 2011-2012 with no
changes from the previous year. Phillip E. abstained from voting.
11. Open Discussion- None.
12. Action items were reviewed and agreed upon.
13. In kind hours were submitted.
14. E-board meeting was scheduled for Friday 7/22/11 at Noon to take place at the LATAG office.
15. The next regular board meeting will be Tuesday August, 9th 2011 at the College.
16. The meeting was ended with unanimous approval at approximately 8:25PM